SupertardMC:

Which equation best represents the data indicated? (graph in comments) A. w = 15c - 75 B. w = 15c - (8/3) C. w = (3/2)c - 75 D. w = (1/14)c - 3 E. w = -10c +170

SupertardMC:

I have attached the graph:

1 attachment
mhanifa:

You need approximate coordinates of the two points to work out the formula. I suggest points (40, 1) and (140, 8)

mhanifa:

Can you find the slope?

mhanifa:

Use slope formula: \[m = \frac{ y_2-y_1 }{ x_2-x_1 }\]

mhanifa:

Are you there?

SupertardMC:

Yes, and I figured it out. It's C. w = (3/2)c - 75 Thanks for helping @mhanifa

mhanifa:

Are you sure C is correct option?

SupertardMC:

Yes. A and B, when graphed, are almost the same (both almost completely vertical) and D and E go in the wrong direction when graphed. C very close

SupertardMC:

To anyone looking at this in the future: Answer C was incorrect.

mhanifa:

The slope is: \[m = \frac{ 8-1 }{ 140-40 } = \frac{ 7 }{ 100 } \approx \frac{ 1 }{ 14}\]

mhanifa:

It was leading to option D without even finding the y-intercept.